PHOTODISSOCIATION OF FORMIC ACID ISOLATED IN SOLID PARAHYDROGEN

\underline{DAVID T. ANDERSON}, LEIF O. PAULSON, Department of Chemistry, University of Wyoming, Laramie, WY 82071-3838.

The in situ photochemistry of dopant molecules isolated in solid parahydrogen (pH_{2}) typically differs from analogous studies in rare gas crystals for two main reasons: (1) solid pH_{2} has a negligible cage effect so that photodissociation of a precursor molecule can lead efficiently to well-separated fragments, and (2) radical fragments can potentially react with the pH_{2} matrix. Our group is currently studying the 193 nm photochemistry of a number of precursor molecules isolated in solid pH_{2} via high-resolution FTIR spectroscopy. In this talk I will present results for the 193 nm photolysis of formic acid (FA) in solid pH_{2}. In rare gas matrices, the analogous FA photochemistry proceeds via the CO+H_{2}O and CO_{2}+H_{2} photodissociation channels.^{a} In solid pH_{2}, in addition to these channels we observe the production of HCO and HOCO. Further, after the UV laser is turned off, the HOCO concentration continues to increase with a slow first-order rate constant for a period of 10 hours. At this point, we still do not have a full explanation of the chemical mechanism leading to the post-photolysis increase in the HOCO concentration.

aJ. Lundell, M. Räsänen, J. Mol. Struct. 436-437, 349 (1997).